milana:

Select the number sentence that illustrates the commutative property of addition. A. 22 + 11 = 33 B. 22 − 11 = 11 C. 22 + 11 = 11 + 22 D. 22 + 11 = 22 + 5 + 6

Extrinix:

Ok so, look for which one adds by it’s opposite basically.

Extrinix:

E.g \(a+b=b+a\)

milana:

um just tell the answer??

Extrinix:

I’m not allowed to do that, but I gave you an easy way to the answer.

milana:

your allowed to tell the answer

Extrinix:

Not by QC rules, just look at the property and what it does. \(a+b=b+a\)

Extrinix:

Can you determine an answer from that?

milana:

can u tell the answer thats all????!!!!

Extrinix:

Just look at your answer choices and find one similar to this. \(a+b=b+a\)

milana:

ok whatever